\begin{thebibliography}{45}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Agirre et~al.(2014)Agirre, de~Lacalle, and Soroa}]{Agirreetal2014}
Eneko Agirre, Oier~Lopez de~Lacalle, and Aitor Soroa. 2014.
\newblock Random walks for knowledge-based word sense disambiguation.
\newblock \emph{Computational Linguistics}, 40(1):57--84.

\bibitem[{Banerjee and Baral(2020)}]{banerjee2020self}
Pratyay Banerjee and Chitta Baral. 2020.
\newblock Self-supervised knowledge triplet learning for zero-shot question
  answering.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 151--162.

\bibitem[{Boella and Di~Caro(2013)}]{BoellaandDiCaro2013}
Guido Boella and Luigi Di~Caro. 2013.
\newblock Supervised learning of syntactic contexts for uncovering definitions
  and extracting hypernym relations in text databases.
\newblock In \emph{Machine learning and knowledge discovery in databases},
  pages 64--79. Springer.

\bibitem[{Bordea et~al.(2015)Bordea, Buitelaar, Faralli, and
  Navigli}]{task17semeval2015}
Georgeta Bordea, Paul Buitelaar, Stefano Faralli, and Roberto Navigli. 2015.
\newblock Semeval-2015 task 17: Taxonomy extraction evaluation (texeval).
\newblock In \emph{Proceedings of the 9th International Workshop on Semantic
  Evaluation}. Association for Computational Linguistics.

\bibitem[{Bordea et~al.(2016)Bordea, Lefever, and
  Buitelaar}]{task13semeval2016}
Georgeta Bordea, Els Lefever, and Paul Buitelaar. 2016.
\newblock Semeval-2016 task 13: Taxonomy extraction evaluation (texeval-2).
\newblock In \emph{Proceedings of the 10th International Workshop on Semantic
  Evaluation}. Association for Computational Linguistics.

\bibitem[{Bouraoui et~al.(2020)Bouraoui, Camacho-Collados, and
  Schockaert}]{bouraoui2020inducing}
Zied Bouraoui, Jose Camacho-Collados, and Steven Schockaert. 2020.
\newblock Inducing relational knowledge from bert.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}, volume~34, pages 7456--7463.

\bibitem[{Chen et~al.(2021)Chen, Lin, and Klein}]{Chen2021ConstructingTF}
Catherine Chen, Kevin Lin, and D.~Klein. 2021.
\newblock Constructing taxonomies from pretrained language models.
\newblock In \emph{NAACL}.

\bibitem[{Chen et~al.(2020)Chen, Chu, Stratos, and Gimpel}]{chen2020mining}
Mingda Chen, Zewei Chu, Karl Stratos, and Kevin Gimpel. 2020.
\newblock Mining knowledge for natural language inference from wikipedia
  categories.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing: Findings}, pages 3500--3511.

\bibitem[{Chiang et~al.(2020)Chiang, Camacho-Collados, and
  Pardos}]{chiang2020understanding}
Hsiao-Yu Chiang, Jose Camacho-Collados, and Zachary Pardos. 2020.
\newblock Understanding the source of semantic regularities in word embeddings.
\newblock In \emph{Proceedings of the 24th Conference on Computational Natural
  Language Learning}, pages 119--131.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{Devlin2019BERTPO}
J.~Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock Bert: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock In \emph{NAACL}.

\bibitem[{Espinosa-Anke et~al.(2021)Espinosa-Anke, Codina-Filb{\'a}, and
  Wanner}]{anke2021evaluating}
Luis Espinosa-Anke, Joan Codina-Filb{\'a}, and Leo Wanner. 2021.
\newblock Evaluating language models for the retrieval and categorization of
  lexical collocations.
\newblock In \emph{Proceedings of the 16th Conference of the European Chapter
  of the Association for Computational Linguistics: Main Volume}, pages
  1406--1417.

\bibitem[{Espinosa-Anke et~al.(2016)Espinosa-Anke, Saggion, Ronzano, and
  Navigli}]{Espinosa-Ankeetal2016AAAI}
Luis Espinosa-Anke, Horacio Saggion, Francesco Ronzano, and Roberto Navigli.
  2016.
\newblock Extasem! extending, taxonomizing and semantifying domain
  terminologies.
\newblock In \emph{Proceedings of AAAI}, Phoenix, USA.

\bibitem[{Fountain and Lapata(2012)}]{FountainandLapata2012}
Trevor Fountain and Mirella Lapata. 2012.
\newblock Taxonomy induction using hierarchical random graphs.
\newblock In \emph{Proceedings of NAACL}, pages 466--476. Association for
  Computational Linguistics.

\bibitem[{Fu et~al.(2014)Fu, Guo, Qin, Che, Wang, and Liu}]{Fuetal2014}
Ruiji Fu, Jiang Guo, Bing Qin, Wanxiang Che, Haifeng Wang, and Ting Liu. 2014.
\newblock Learning semantic hierarchies via word embeddings.
\newblock In \emph{Proceedings of ACL}, volume~1.

\bibitem[{Garcia et~al.(2021)Garcia, Vieira, Scarton, Idiart, and
  Villavicencio}]{garcia2021probing}
Marcos Garcia, Tiago~Kramer Vieira, Carolina Scarton, Marco Idiart, and Aline
  Villavicencio. 2021.
\newblock Probing for idiomaticity in vector space models.
\newblock In \emph{Proceedings of the 16th Conference of the European Chapter
  of the Association for Computational Linguistics: Main Volume}, pages
  3551--3564.

\bibitem[{Grefenstette(2015)}]{grefenstette2015inriasac}
Gregory Grefenstette. 2015.
\newblock Inriasac: Simple hypernym extraction methods.
\newblock In \emph{Proceedings of the 9th International Workshop on Semantic
  Evaluation (SemEval 2015)}, pages 911--914.

\bibitem[{Hearst(1992)}]{Hearst1992}
Marti~A Hearst. 1992.
\newblock Automatic acquisition of hyponyms from large text corpora.
\newblock In \emph{Proceedings of the 14th conference on Computational
  linguistics}, pages 539--545.

\bibitem[{Hwang et~al.(2012)Hwang, Grauman, and Sha}]{Hwangetal2012}
Sung~Ju Hwang, Kristen Grauman, and Fei Sha. 2012.
\newblock Semantic kernel forests from multiple taxonomies.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  1718--1726.

\bibitem[{Jiang et~al.(2020)Jiang, Xu, Araki, and Neubig}]{Jiang2020HowCW}
Zhengbao Jiang, Frank~F. Xu, J.~Araki, and Graham Neubig. 2020.
\newblock How can we know what language models know?
\newblock \emph{Transactions of the Association for Computational Linguistics},
  8:423--438.

\bibitem[{Joshi et~al.(2020)Joshi, Lee, Luan, and
  Toutanova}]{joshi2020contextualized}
Mandar Joshi, Kenton Lee, Yi~Luan, and Kristina Toutanova. 2020.
\newblock Contextualized representations using textual encyclopedic knowledge.
\newblock \emph{arXiv preprint arXiv:2004.12006}.

\bibitem[{Kassner et~al.(2021)Kassner, Dufter, and
  Sch{\"u}tze}]{kassner-etal-2021-multilingual}
Nora Kassner, Philipp Dufter, and Hinrich Sch{\"u}tze. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.eacl-main.284} {Multilingual
  {LAMA}: Investigating knowledge in multilingual pretrained language models}.
\newblock In \emph{Proceedings of the 16th Conference of the European Chapter
  of the Association for Computational Linguistics: Main Volume}, pages
  3250--3258, Online. Association for Computational Linguistics.

\bibitem[{Kozareva and Hovy(2010)}]{KozarevaandHovy2010}
Zornitsa Kozareva and Eduard Hovy. 2010.
\newblock A semi-supervised method to learn and construct taxonomies using the
  web.
\newblock In \emph{Proceedings of EMNLP}, pages 1110--1118.

\bibitem[{Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis,
  Zettlemoyer, and Stoyanov}]{Liu2019RoBERTaAR}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, M.~Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019.
\newblock Roberta: A robustly optimized bert pretraining approach.
\newblock \emph{ArXiv}, abs/1907.11692.

\bibitem[{Mao et~al.(2018)Mao, Ren, Shen, Gu, and Han}]{Mao2018EndtoEndRL}
Yuning Mao, Xiang Ren, J.~Shen, X.~Gu, and Jiawei Han. 2018.
\newblock End-to-end reinforcement learning for automatic taxonomy induction.
\newblock In \emph{ACL}.

\bibitem[{Panchenko et~al.(2016)Panchenko, Faralli, Ruppert, Remus, Naets,
  Fairon, Ponzetto, and Biemann}]{Panchenko2016TAXIAS}
Alexander Panchenko, Stefano Faralli, E.~Ruppert, Steffen Remus, Hubert Naets,
  Cedric Fairon, Simone~Paolo Ponzetto, and Chris Biemann. 2016.
\newblock Taxi at semeval-2016 task 13: a taxonomy induction method based on
  lexico-syntactic patterns, substrings and focused crawling.
\newblock In \emph{SemEval@NAACL-HLT}.

\bibitem[{Paranjape et~al.(2021)Paranjape, Michael, Ghazvininejad, Zettlemoyer,
  and Hajishirzi}]{paranjape2021prompting}
Bhargavi Paranjape, Julian Michael, Marjan Ghazvininejad, Luke Zettlemoyer, and
  Hannaneh Hajishirzi. 2021.
\newblock Prompting contrastive explanations for commonsense reasoning tasks.
\newblock \emph{arXiv preprint arXiv:2106.06823}.

\bibitem[{Petroni et~al.(2019)Petroni, Rockt{\"a}schel, Lewis, Bakhtin, Wu,
  Miller, and Riedel}]{Petroni2019LanguageMA}
Fabio Petroni, Tim Rockt{\"a}schel, Patrick Lewis, A.~Bakhtin, Yuxiang Wu,
  Alexander~H. Miller, and S.~Riedel. 2019.
\newblock Language models as knowledge bases?
\newblock In \emph{EMNLP}.

\bibitem[{Pilehvar et~al.(2013)Pilehvar, Jurgens, and
  Navigli}]{Pilehvaretal:2013}
Mohammad~Taher Pilehvar, David Jurgens, and Roberto Navigli. 2013.
\newblock {A}lign, {D}isambiguate and {W}alk: a {U}nified {A}pproach for
  {M}easuring {S}emantic {S}imilarity.
\newblock In \emph{Proceedings of ACL}, pages 1341--1351, Sofia, Bulgaria.

\bibitem[{Qin and Eisner(2021)}]{qin2021learning}
Guanghui Qin and Jason Eisner. 2021.
\newblock Learning how to ask: Querying lms with mixtures of soft prompts.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 5203--5212.

\bibitem[{Radford et~al.(2019)Radford, Wu, Child, Luan, Amodei, and
  Sutskever}]{radford2019language}
Alec Radford, Jeff Wu, Rewon Child, David Luan, Dario Amodei, and Ilya
  Sutskever. 2019.
\newblock Language models are unsupervised multitask learners.

\bibitem[{Sainz and Rigau(2021)}]{sainz2021ask2transformers}
Oscar Sainz and German Rigau. 2021.
\newblock Ask2transformers: Zero-shot domain labelling with pretrained language
  models.
\newblock In \emph{Proceedings of the 11th Global Wordnet Conference}, pages
  44--52.

\bibitem[{Salazar et~al.(2020)Salazar, Liang, Nguyen, and
  Kirchhoff}]{Salazar2020MaskedLM}
Julian Salazar, Davis Liang, Toan~Q. Nguyen, and Katrin Kirchhoff. 2020.
\newblock Masked language model scoring.
\newblock In \emph{ACL}.

\bibitem[{Shang et~al.(2020)Shang, Dash, Chowdhury, Mihindukulasooriya, and
  Gliozzo}]{Shang2020TaxonomyCO}
Chao Shang, Sarthak Dash, Md. Faisal~Mahbub Chowdhury, Nandana
  Mihindukulasooriya, and A.~Gliozzo. 2020.
\newblock Taxonomy construction of unseen domains via graph-based cross-domain
  knowledge transfer.
\newblock In \emph{ACL}.

\bibitem[{Shin et~al.(2020)Shin, Razeghi, Logan~IV, Wallace, and
  Singh}]{shin2020eliciting}
Taylor Shin, Yasaman Razeghi, Robert~L Logan~IV, Eric Wallace, and Sameer
  Singh. 2020.
\newblock Eliciting knowledge from language models using automatically
  generated prompts.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 4222--4235.

\bibitem[{Shwartz et~al.(2020)Shwartz, West, Le~Bras, Bhagavatula, and
  Choi}]{shwartz2020unsupervised}
Vered Shwartz, Peter West, Ronan Le~Bras, Chandra Bhagavatula, and Yejin Choi.
  2020.
\newblock Unsupervised commonsense question answering with self-talk.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 4615--4629.

\bibitem[{Snow et~al.(2004)Snow, Jurafsky, and Ng}]{Snowetal2004}
Rion Snow, Daniel Jurafsky, and Andrew~Y Ng. 2004.
\newblock Learning syntactic patterns for automatic hypernym discovery.
\newblock \emph{Advances in Neural Information Processing Systems 17}.

\bibitem[{Ushio et~al.(2021)Ushio, Espinosa-Anke, Schockaert, and
  Camacho-Collados}]{ushio-etal-2021-bert-is}
Asahi Ushio, Luis Espinosa-Anke, Steven Schockaert, and Jose Camacho-Collados.
  2021.
\newblock {BERT} is to {NLP} what {A}lex{N}et is to {CV}: {C}an {P}re-{T}rained
  {L}anguage {M}odels {I}dentify {A}nalogies?
\newblock In \emph{Proceedings of the {ACL}-{IJCNLP} 2021 Main Conference}.
  Association for Computational Linguistics.

\bibitem[{Velardi et~al.(2013)Velardi, Faralli, and Navigli}]{Velardietal2013}
Paola Velardi, Stefano Faralli, and Roberto Navigli. 2013.
\newblock {OntoLearn Reloaded}: A graph-based algorithm for taxonomy induction.
\newblock \emph{Computational Linguistics}, 39(3):665--707.

\bibitem[{Wang and Cho(2019)}]{wang-cho-2019-bert}
Alex Wang and Kyunghyun Cho. 2019.
\newblock \href {https://doi.org/10.18653/v1/W19-2304} {{BERT} has a mouth, and
  it must speak: {BERT} as a {M}arkov random field language model}.
\newblock In \emph{Proceedings of the Workshop on Methods for Optimizing and
  Evaluating Neural Language Generation}, pages 30--36, Minneapolis, Minnesota.
  Association for Computational Linguistics.

\bibitem[{Wolf et~al.(2020)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac,
  Rault, Louf, Funtowicz, Davison, Shleifer, von Platen, Ma, Jernite, Plu, Xu,
  Scao, Gugger, Drame, Lhoest, and Rush}]{wolf-etal-2020-transformers}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue,
  Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, Joe
  Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien
  Plu, Canwen Xu, Teven~Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest,
  and Alexander~M. Rush. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.emnlp-demos.6}
  {Transformers: State-of-the-art natural language processing}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing: System Demonstrations}, pages 38--45, Online.
  Association for Computational Linguistics.

\bibitem[{Yang and Callan(2009)}]{YangandCallan2009}
Hui Yang and Jamie Callan. 2009.
\newblock A metric-based framework for automatic taxonomy induction.
\newblock In \emph{Proceedings of ACL/IJCNLP}, pages 271--279. Association for
  Computational Linguistics.

\bibitem[{Yu and Dredze(2014)}]{yu2014improving}
Mo~Yu and Mark Dredze. 2014.
\newblock Improving lexical embeddings with semantic knowledge.
\newblock In \emph{ACL (2)}, pages 545--550.

\bibitem[{Yu et~al.(2020)Yu, Li, Shen, Feng, Sun, and Zhang}]{Yu2020STEAMST}
Yue Yu, Yinghao Li, Jiaming Shen, Haoyang Feng, Jimeng Sun, and Chao Zhang.
  2020.
\newblock Steam: Self-supervised taxonomy expansion with mini-paths.
\newblock \emph{Proceedings of the 26th ACM SIGKDD International Conference on
  Knowledge Discovery \& Data Mining}.

\bibitem[{Yu et~al.(2015)Yu, Wang, Lin, and Wang}]{Yuetal2015}
Zheng Yu, Haixun Wang, Xuemin Lin, and Min Wang. 2015.
\newblock Learning term embeddings for hypernymy identification.
\newblock In \emph{Proceedings of IJCAI}, pages 1390--1397.

\bibitem[{Zhou et~al.(2019)Zhou, Ge, Xu, Wei, and Zhou}]{zhou2019bert}
Wangchunshu Zhou, Tao Ge, Ke~Xu, Furu Wei, and Ming Zhou. 2019.
\newblock Bert-based lexical substitution.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 3368--3373.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{acl}

\usepackage{url}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{comment}

\usepackage{microtype}

\usepackage{amsmath, amssymb}
\usepackage{xcolor}
\usepackage{todonotes}
\usepackage{booktabs}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{float}
\usepackage{subcaption}

\newcommand{\luis}[1]{\textcolor{blue}{#1}}
\newcommand{\devansh}[1]{\textcolor{red}{#1}}
\newcommand{\todoin}[1]{\todo[inline,size=\small,color=red!10]{\textbf{all: }#1}\xspace}

\DeclareMathOperator*{\argmax}{arg\,max}

\title{Distilling Hypernymy Relations from Language Models: On the Effectiveness of Zero-Shot Taxonomy Induction}

\author{\textbf{Devansh Jain$^{\spadesuit}$, Luis Espinosa Anke$^{\diamondsuit}$}\\
$^{\spadesuit}$ Department of Computer Science and Information Systems, BITS Pilani, India \\
  $^{\diamondsuit}$ Cardiff NLP, School of Computer Science and Informatics, Cardiff University, UK  \\
  $^{\spadesuit}$ \texttt{f20180798@pilani.bits-pilani.ac.in} ,
   $^{\diamondsuit}$ \texttt{espinosa-ankel@cardiff.ac.uk} 
 }  

\begin{document}
\maketitle
\begin{abstract}
In this paper, we analyze zero-shot taxonomy learning methods which are based on distilling knowledge from language models via prompting and sentence scoring. We show that, despite their simplicity, these methods outperform some supervised strategies and are competitive with the current state-of-the-art under adequate conditions. We also show that statistical and linguistic properties of prompts dictate downstream performance.
\end{abstract}

\section{Introduction}
\label{sec:intro}

Taxonomy learning (TL) is the task of arranging domain terminologies into hierarchical structures where terms are nodes  and edges denote \textit{is-a} (hypernymic) relationships \cite{Hwangetal2012}. Domain-specific concept generalization is at the core of human cognition \cite{Yuetal2015}, and a key enabler in NLP tasks where inference and reasoning are important, e.g.: semantic similarity \cite{Pilehvaretal:2013, yu2014improving}, WSD \cite{Agirreetal2014} and, more recently, QA \cite{joshi2020contextualized} and NLI \cite{chen2020mining}.

Earlier approaches to taxonomy learning focused on mining lexico-syntactic patterns from candidate $($hyponym, hypernym$)$ pairs \cite{Hearst1992,Snowetal2004,KozarevaandHovy2010,BoellaandDiCaro2013,Espinosa-Ankeetal2016AAAI}, clustering \cite{YangandCallan2009}, graph-based methods \cite{FountainandLapata2012,Velardietal2013} or word embeddings \cite{Fuetal2014,Yuetal2015}. These methods, which largely rely on hand-crafted features, are still relevant today, and complement modern approaches exploiting language models (LMs), either via sequence classification \cite{Chen2021ConstructingTF}, or combining contextual, distributed, and lexico-syntactic features \cite{Yu2020STEAMST}. In parallel, several works have recently focused on using LMs as zero-shot tools for solving NLP tasks, e.g., commonsense, relational and analogical reasoning \cite{Petroni2019LanguageMA,bouraoui2020inducing, ushio-etal-2021-bert-is,paranjape2021prompting}, multiword expression (MWE) identification \cite{anke2021evaluating,garcia2021probing}, QA \cite{shwartz2020unsupervised, banerjee2020self}, domain labeling \cite{sainz2021ask2transformers}, or lexical substitution and simplification \cite{zhou2019bert}. Moreover, by tuning and manipulating natural language queries (often referred to as \textit{prompts}),
impressive results have been recently obtained on tasks such as semantic textual similarity, entailment, or relation classification \cite{shin2020eliciting,qin2021learning}.

In this paper, we evaluate LMs on TL benchmarks using prompt-based and sentence-scoring techniques, and find not only that they are competitive with common approaches proposed in the literature (which are typically supervised and/or reliant on external resources), but that they achieve SoTa results in certain domains.

\section{Methodology}
\label{sec:methodology}
We follow \citet{ushio-etal-2021-bert-is} and define a prompt generation function $\tau_p(t_1, t_2)$ which maps a pair of terms and a prompt type $p$ to a single sentence. For instance,
\begin{equation*}
    \begin{aligned}
        \tau_{kind}\text{(``physics'', ``science'')} = \\
        \text{``physics is a kind of science''}
    \end{aligned}
\end{equation*}

\noindent Then, given a terminology $\mathcal{T}$, the goal is to, given an input term $t \in \mathcal{T}$, retrieve its top $k$ most likely hypernyms, (in our experiments, $k \in \{1, 3, 5\}$), using either masked language model (MLM) prompting ($\S$\ref{subsec:restrict}), or sentence-scoring ($\S$\ref{subsec:scorer}).

\subsection{MLM Prompting}
\label{subsec:restrict}

\paragraph{RestrictMLM} \citet{Petroni2019LanguageMA} introduced a ``fill-in-the-blanks'' approach based on cloze statements (or \textit{prompts}) to extract relational knowledge from pretrained LMs. The intuition being that an LM can be considered to ``know'' a fact (in the form of a $<$\emph{subject, relation, object}$>$ triple) such as $<$\emph{Madrid, capital-of, Spain}$>$ if it can successfully predict the correct words when queried with prompts such as ``Madrid is the capital of {\texttt{[MASK]}}''. 
We extend this formulation to define a hypernym retrieval function $f_{R}(\cdot)$ as follows:
\begin{equation}
    \begin{aligned}
        f_{R}(p, t, \textbf{T}) = P({\small{\texttt{[MASK]}}} | \tau_p(t, {\small{\texttt{[MASK]}}})) * \textbf{T}
    \end{aligned}
\end{equation}
where $p$ is a prompt type, and $\textbf{T}$ is a one-hot encoding of the terms $\mathcal{T}$ in the LM's vocabulary. We follow previous works \citep{Petroni2019LanguageMA, kassner-etal-2021-multilingual} and restrict the output probability distribution since this task requires the construction of a lexical taxonomy starting from a fixed vocabulary.

\paragraph{PromptMLM} For completeness, we also report results for an unrestricted variant of \textit{RestrictMLM}, where the LM's entire vocabulary is considered.

\subsection{LMScorer}
\label{subsec:scorer}
Factual (and true) information such as ``Trout is a type of fish'' should be scored higher by a LM than fictitious information such as ``Trout is a type of mammal''. The method for scoring a sentence depends on the type of LM used.

\paragraph{Causal Language Models} Given a sentence $\mathbf{W}$, causal LMs ($\mathcal{C}$) predict token $w_i$ using only past tokens $\mathbf{W}_{<i}$. Thus, a likelihood score can be estimated for each token $w_i$ from the LM's next token prediction. The corresponding scores are then aggregated to yield a score for the sentence $\mathbf{W}$.
\begin{equation}
    s_{\mathcal{C}}(\mathbf{W}) = \exp \left(\sum_{i=1}^{|\mathbf{W}|} log P_{\mathcal{C}}(w_i | \mathbf{W}_{<i})\right)
\end{equation}

\paragraph{Masked Language Models}
Given a sentence $\mathbf{W}$, masked LMs ($\mathcal{M}$) replace $w_i$ by \texttt{[MASK]} and predict it using past and future tokens. Thus, a pseudo-likelihood score can be computed for each token $w_i$ by iteratively masking it and using the LM's masked token prediction \cite{wang-cho-2019-bert, Salazar2020MaskedLM}. The corresponding scores are then aggregated to yield a score for the sentence $\mathbf{W}$.

\begin{equation}
    s_{\mathcal{M}}(\mathbf{W}) = \exp \left(\sum_{i=1}^{|\mathbf{W}|} log P_{\mathcal{M}}(w_i | \mathbf{W}_{\setminus i})\right)
\end{equation}

\noindent Given the above, we can cast TL as a sentence-scoring problem by evaluating the natural fluency of hypernymy-eliciting sentences. Specifically, for each term $t$, we score the sentences generated using $\tau_p(\cdot)$ with every other term $t'$ in the terminology. We then select the term-pair with the highest sentence score and assume that the corresponding term $t'$ is a hypernym of $t$. Formally, we define a hypernym selection function $f_S(\cdot)$ as follows:
 \begin{equation}
    f_{S}(p, t, \mathcal{T}) = \argmax_{t' \in \mathcal{T} \setminus t} [s(\tau_p(t, t'))]
\end{equation}
where $s$ refers to the scoring function determined by the LM used.

\section{Experimental setup}
\label{sec:experiments}

This section covers the datasets and prompts we use in our experiments\footnote{We use PyTorch and the {\small{\texttt{transformers}}} library \cite{wolf-etal-2020-transformers}, as well as {\small{\texttt{mlm-scoring}}} \cite{Salazar2020MaskedLM} (\url{https://github.com/awslabs/mlm-scoring}).}, as well as the different LMs we consider. Concerning evaluation metrics, we report standard precision ($P$), recall ($R$) and $F$-score at the \textit{edge level} \cite{task13semeval2016}.

\paragraph{Dataset Details} We evaluate our proposed approaches on datasets belonging to two TL SemEval tasks (\textit{TExEval-1}, \citet{task17semeval2015} and \textit{TExEval-2}, \citet{task13semeval2016}). Following recent literature, we consider the \textit{equipment} taxonomy from \textit{TExEval-1} and the English-language \textit{environment}, \textit{science} and \textit{food} taxonomies from \textit{TExEval-2}. For the \textit{science} taxonomy, our results are based on an \textit{average of the 3 subsets}, which is in line with previous work. Since these datasets do not come with training data, they are well suited for unsupervised approaches.

\begin{center}
\begin{table}[h]
\centering
\small
\begin{tabular}{@{}l|l|r|r@{}}
\toprule
Domain & Source & \multicolumn{1}{c|}{$V$} & \multicolumn{1}{c}{$E$} \\ \midrule
\textit{environment} & Eurovoc & 261 & 261 \\ \midrule
\multirow{3}{*}{\textit{science}} & Combined & 453 & 465 \\
 & Eurovoc & 125 & 124 \\
 & WordNet & 429 & 452 \\ \midrule
\textit{food} & Combined & 1556 & 1587 \\ \midrule
\textit{equipment} & Combined & 612 & 615 \\ \bottomrule
\end{tabular}
\caption{Taxonomies statistics. Vertices ($V$) and Edges ($E$) are often used as structural measures.}
\label{tab:dataset}
\end{table}
\end{center}

\paragraph{Prompts} We use the following prompts:
\begin{itemize}
    \item \textit{gen.}: [$t_2$] is more general than [$t_1$].
    \vspace{-0.8ex} \item \textit{spec.}: [$t_1$] is more specific than [$t_2$].
    \vspace{-0.8ex} \item \textit{type}: [$t_1$] is a type of [$t_2$].
\end{itemize}
$gen.$ and $spec.$ prompts are hand-crafted templates to encode, in a general way, the hypernymy relationship. The choice of the $type$ prompt, however, comes from a set of experiments involving all \textit{LPAQA} \cite{Jiang2020HowCW} prompts under the ``\textit{is  a  subclass  of}'' category. We do not consider automatic prompt generation techniques \cite{shin2020eliciting} due to the absence of training data.
Note that for each prompt, we replace $t_1$ with the input term so that the task is always to predict its hypernym.

\paragraph{Language Models} We interrogate BERT \cite{Devlin2019BERTPO} and RoBERTa \cite{Liu2019RoBERTaAR} among masked LMs, and GPT2 \cite{radford2019language} among causal LMs. For each LM, we consider two variants corresponding to approximately 117M parameters and 345M parameters.

\section{Results}
\label{sec:results}

Table~\ref{tab:sci-ev} shows the results on \textit{TExEval-2}'s \textit{science} and \textit{environment}. We compare with the current state of the art (\textit{Graph2Taxo}) \cite{Shang2020TaxonomyCO}, as well as with other strong baselines such as \textit{TaxoRL} \cite{Mao2018EndtoEndRL} and \textit{TAXI} \cite{Panchenko2016TAXIAS}, the highest ranked system in \textit{TExEval-2}. We also compare with \textit{CTP} \cite{Chen2021ConstructingTF} to illustrate the advantages of zero-shot methods vs finetuning. For the \textit{environment} domain, we find that \textit{RestrictMLM} performs similar to \textit{CTP} and  \textit{LMScorer} outperforms it. Moreover, all 3 proposed approaches fail to outperform the other baselines. However, in \textit{science}, all 3 of our approaches outperform \textit{CTP}, while our best model (\textit{RestrictMLM}) outperforms \textit{TAXI} and is competitive with \textit{TaxoRL} (ours has higher precison, but lower recall). Note that compared to our zero-shot approaches, these methods are either supervised, expensive to train or take advantage of external taxonomical resources such as WordNet, or lexico-syntactic patterns mined from the web using different hand-crafted heuristics.

\begin{table}[h]
\resizebox{\columnwidth}{!}{
\begin{tabular}{@{}l|rrr|rrr@{}}
\toprule
                              & \multicolumn{3}{c|}{\textit{environment}}                                       & \multicolumn{3}{c}{\textit{science}}                                      \\ \midrule
Model                                & \multicolumn{1}{c}{P} & \multicolumn{1}{c}{R} & \multicolumn{1}{c|}{F} & \multicolumn{1}{c}{P} & \multicolumn{1}{c}{R} & \multicolumn{1}{c}{F} \\ \midrule
\textit{TAXI}       & 33.8                  & 26.8                  & 29.9                   & 35.2                  & 35.3                  & 35.2                  \\
\textit{TaxoRL}     & 32.3                  & 32.3                  & 32.3                   & 37.9                  & 37.9                  & 37.9                  \\
\textit{Graph2Taxo} & 89.0                  & 24.0                  & \textbf{37.0}                   & 84.0                  & 30.0                  & \textbf{44.0}                  \\
\textit{CTP}        & 23.1                  & 23.0                  & 23.0                   & 29.4                  & 28.8                  & 29.1                  \\ \midrule
\textit{PromptMLM}                        & 19.2                  & 19.2                  & 19.2                   & 34.4                  & 32.0                  & 33.1                  \\
\textit{RestrictMLM}                        & 23.0                  & 23.0                  & 23.0                   & 39.3                  & 36.7                  & 37.9                  \\
\textit{LMScorer}                         & 26.4                  & 26.4                  & 26.4                   & 33.1                  & 30.7                  & 31.8                  \\ \bottomrule
\end{tabular}
}
\caption{Comparison of our best performing methods with previous work (\textit{environment} and \textit{science}).}
\label{tab:sci-ev}
\end{table}

We also show results for \textit{TExEval-1}'s \textit{equipment} and \textit{TExEval-2}'s \textit{food} (Table~\ref{tab:food-eq}). Both datasets are considerably larger than \textit{environment} and \textit{science}. We compare with the corresponding highest ranked system, namely \textit{TAXI} for \textit{food}, and \textit{INRIASAC} \cite{grefenstette2015inriasac} for \textit{equipment}. For both domains, all 3 of our approaches outperform the corresponding \textit{TExEval} best-performing systems. This suggests that zero-shot TL with LMs is robust, easily scalable and feasible on large taxonomies.

Finally, a clear bottleneck for prompt-based methods is that only single-token terms can be predicted (using a single {\small{\texttt{[MASK]}}} token), making this approach a lower bound for TL.

\begin{table}[h]
\resizebox{\columnwidth}{!}{
\begin{tabular}{@{}l|rrr|rrr@{}}
\toprule
                           & \multicolumn{3}{c|}{\textit{food}}                                           & \multicolumn{3}{c}{\textit{equipment}}                                   \\ \midrule
Model                             & \multicolumn{1}{c}{P}    & \multicolumn{1}{c}{R} & \multicolumn{1}{c|}{F} & \multicolumn{1}{c}{P} & \multicolumn{1}{c}{R} & \multicolumn{1}{c}{F} \\ \midrule
\textit{TExEval} & \multicolumn{1}{l}{13.2} & 25.1                  & 17.3                   & 51.8                  & 18.8                  & 27.6                  \\ \midrule
\textit{PromptMLM}                     & 23.2                     & 22.6                  & 22.9                   & 29.4                  & 29.3                  & 29.4                  \\
\textit{RestrictMLM}                     & 25.2                     & 24.6                  & \textbf{24.9}                   & 38.4                  & 38.2                  & \textbf{38.3}                  \\
\textit{LMScorer}                    & 25.2                     & 24.6                  & \textbf{24.9}                   & 37.7                  & 37.6                  & 37.6                  \\ \bottomrule
\end{tabular}
}
\caption{Comparison of our best configurations with the best TExEval systems on \textit{food} and \textit{equipment}.}
\label{tab:food-eq}
\end{table}

\begin{table*}[!th]
\resizebox{\textwidth}{!}{
\begin{tabular}{@{}l|l|c|rrr|c|rrr|c|rrr|crrr@{}}
\toprule
 &  & \multicolumn{4}{c|}{\textit{environment}} & \multicolumn{4}{c|}{\textit{science}} & \multicolumn{4}{c|}{\textit{food}} & \multicolumn{4}{c}{\textit{equipment}} \\ \midrule
Method & LM & $(p, k)$ & \multicolumn{1}{c}{P} & \multicolumn{1}{c}{R} & \multicolumn{1}{c|}{F} & $(p, k)$ & \multicolumn{1}{c}{P} & \multicolumn{1}{c}{R} & \multicolumn{1}{c|}{F} & $(p, k)$ & \multicolumn{1}{c}{P} & \multicolumn{1}{c}{R} & \multicolumn{1}{c|}{F} & \multicolumn{1}{c|}{$(p, k)$} & \multicolumn{1}{c}{P} & \multicolumn{1}{c}{R} & \multicolumn{1}{c}{F} \\ \midrule
\multirow{4}{*}{\textit{PromptMLM}} & BERT-Base & $(t, 1)$ & 18.8 & 18.8 & 18.8 & $(t, 1)$ & 30.2 & 28.1 & 29.1 & $(t, 1)$ & 20.9 & 20.4 & 20.6 & \multicolumn{1}{c|}{$(t, 1)$} & 29.4 & 29.3 & 29.4 \\
 & BERT-Large & $(t, 1)$ & 19.2 & 19.2 & 19.2 & $(t, 1)$ & 34.4 & 32.0 & 33.1 & $(t, 1)$ & 23.2 & 22.6 & 22.9 & \multicolumn{1}{c|}{$(t, 1)$} & 28.4 & 28.3 & 28.4 \\
 & RoBERTa-Base & $(t, 1)$ & 18.0 & 18.0 & 18.0 & $(t, 1)$ & 24.5 & 23.0 & 23.7 & $(t, 1)$ & 18.5 & 18.0 & 18.2 & \multicolumn{1}{c|}{$(t, 1)$} & 26.3 & 26.2 & 26.3 \\
 & RoBERTa-Large & $(t, 1)$ & 18.0 & 18.0 & 18.0 & $(t, 1)$ & 28.1 & 26.2 & 27.1 & $(t, 1)$ & 20.3 & 19.8 & 20.0 & \multicolumn{1}{c|}{$(t, 1)$} & 28.4 & 28.3 & 28.4 \\
\midrule
\multirow{4}{*}{\textit{RestrictMLM}} & BERT-Base & $(t, 1)$ & 23.0 & 23.0 & 23.0 & $(t, 1)$ & 35.8 & 33.5 & 34.6 & $(t, 1)$ & 22.8 & 22.2 & 22.5 & \multicolumn{1}{c|}{$(t, 1)$} & 38.4 & 38.2 & 38.3 \\
 & BERT-Large & $(t, 1)$ & 21.8 & 21.8 & 21.8 & $(t, 1)$ & 39.3 & 36.7 & \textbf{37.9} & $(t, 1)$ & 25.2 & 24.6 & \textbf{24.9} & \multicolumn{1}{c|}{$(t, 1)$} & 37.9 & 37.7 & \textbf{37.8} \\
 & RoBERTa-Base & $(t, 1)$ & 5.4 & 5.4 & 5.4 & $(t, 1)$ & 11.0 & 10.6 & 10.8 & $(t, 1)$ & 9.3 & 9.1 & 9.2 & \multicolumn{1}{c|}{$(t, 1)$} & 0.0 & 0.0 & 0.0 \\
 & RoBERTa-Large & $(t, 1)$ & 8.4 & 8.4 & 8.4 & $(t, 1)$ & 12.3 & 11.8 & 12.0 & $(t, 1)$ & 10.7 & 10.5 & 10.6 & \multicolumn{1}{c|}{$(t, 1)$} & 0.0 & 0.0 & 0.0 \\
\midrule
\multirow{6}{*}{\textit{LMScorer}} & BERT-Base & $(t, 1)$ & 20.3 & 20.3 & 20.3 & $(t, 1)$ & 15.2 & 14.4 & 14.8 & $(t, 3)$ & 6.8 & 19.7 & 10.1 & \multicolumn{1}{c|}{$(t, 3)$} & 7.5 & 22.4 & 11.2 \\
 & BERT-Large & $(t, 3)$ & 13.7 & 41.0 & 20.5 & $(t, 1)$ & 13.0 & 12.4 & 12.6 & $(t, 1)$ & 13.9 & 13.6 & 13.7 & \multicolumn{1}{c|}{$(t, 1)$} & 15.2 & 15.1 & 15.1 \\
 & RoBERTa-Base & $(g, 3)$ & 7.7 & 23.0 & 11.5 & $(t, 3)$ & 5.5 & 15.7 & 8.1 & $(t, 3)$ & 2.5 & 7.2 & 3.7 & \multicolumn{1}{c|}{$(t, 5)$} & 4.2 & 21.0 & 7.0 \\
 & RoBERTa-Large & $(t, 3)$ & 11.1 & 33.3 & 16.7 & $(t, 1)$ & 13.6 & 12.8 & 13.2 & $(t, 3)$ & 3.6 & 10.6 & 5.4 & \multicolumn{1}{c|}{$(t, 3)$} & 9.2 & 27.5 & 13.8 \\
 & GPT-2 Base & $(t, 1)$ & 24.9 & 24.9 & 24.9 & $(t, 1)$ & 29.3 & 27.4 & 28.3 & $(t, 1)$ & 21.0 & 20.5 & 20.7 & \multicolumn{1}{c|}{$(t, 1)$} & 36.8 & 36.6 & 36.7 \\
 & GPT-2 Medium & $(t, 1)$ & 26.4 & 26.4 & \textbf{26.4} & $(t, 1)$ & 33.1 & 30.7 & 31.8 & $(t, 1)$ & 25.2 & 24.6 & \textbf{24.9} & \multicolumn{1}{c|}{$(t, 1)$} & 37.7 & 37.6 & 37.7 \\ \bottomrule
\end{tabular}
}
\caption{Comparison of best configuration for each LM and proposed approach. $(p, k)$ refers to the prompt and top-$k$ combination that gives the best results for that setting, where $p$ = $g$ for \textit{gen.}, $s$ for \textit{spec.} and $t$ for \textit{type} prompt.}
\label{tab:lm-compare}
\end{table*}

\section{Analysis}
\label{sec:analysis}

In this section, we provide an in-depth analysis of our approaches, including comparison of LMs and statistical and semantic properties of prompts.

\paragraph{LM Comparison} Table \ref{tab:lm-compare} compares the best configuration for each LM. We can immediately see that a conservative approach (i.e., $k=1$ with the \textit{type} prompt) almost always yields the best $F$-score. Another important conclusion is that, among MLMs, BERT-Large performs best across the board, with BERT generally outperforming RoBERTa, a finding in line with previous works \cite{shin2020eliciting}. Concerning causal LMs, GPT-2 Medium outperforms its smaller counterpart as well as both MLMs for sentence-scoring.

\paragraph{Sensitivity to Prompts} There is interested in understanding the model's sensitivity to prompts and whether frequency can explain downstream performance in lexical semantics tasks \cite{chiang2020understanding}. In the context of prompt vs. performance correlation, we find that prompt-based downstream performance on TL can be attributed to: (1) syntactic completeness and (2) semantic correctness. For (1), we find that prompts that are syntactically more complete (e.g., ``\textit{[X] is a type of [Y]}'' vs ``\textit{[X] is a type [Y]}'', the difference being the prepositional phrase) perform better. For (2), we find that prompts that unambiguously encode hypernymy are also better (i.e., the \textit{type} prompt, as opposed to other noise-inducing templates such as ``\textit{is a}'' or ``\textit{is kind of}''). Finally, out of the cleanest prompts, the most frequent in pretraining corpora are the most competitive. Table \ref{tab:freqs} confirms the intuition that the \textit{type} prompt is not only unambiguous, but also highly frequent when compared to similar (noise-free and syntactically complete) prompts.

\begin{center}
\begin{table}[!h]
\small
\centering
\begin{tabular}{@{}l|r||r@{}}
\toprule
Prompt & \multicolumn{1}{c||}{\textit{avg} F} & \multicolumn{1}{c}{Frequency} \\ \midrule
is a type of & 25.5 & 14,503 \\
is the type of & 24.2 & 809 \\
is a kind of & 23.6 & 2,934 \\
is a form of & 22.1 & 9,518 \\
is one form of & 17.9 & 124 \\
is a & 7.4 & 9,328,426 \\
is a type & 1.0 & 15,085 \\ \bottomrule
\end{tabular}
\caption{Domain-wise average $F$-score of LPAQA prompts and their frequency in BERT's pretraining corpora.}
\label{tab:freqs}
\end{table}
\end{center}

\paragraph{Single-Token vs Multi-Token Hypernyms} 
Table~\ref{tab:single-stats} compares \textit{F-score} on original terminology vs filtered terminology, where filtered terminology contains only the terms that have single-token hypernyms. The results show that \% Increase in \textit{F-score} is inversely proportional to the \% Retained. This can be explained by the fact that smaller \% of terms retained implies higher \% of multi-token hypernyms in the original dataset that cannot be predicted using prompting. Thus, the increase in \textit{F-score} by removing such hypernyms should increase as the \% Retained decreases.

\begin{table}[h]
\resizebox{\columnwidth}{!}{
\begin{tabular}{@{}l|r|r|r@{}}
\toprule
Domain               & \multicolumn{1}{l|}{Total Terms} & \multicolumn{1}{l|}{\% Retained} & \multicolumn{1}{l}{\% Increase} \\ \midrule
\textit{environment} & 261                              & 29.89                            & 2.32                            \\
\textit{equipment}   & 612                              & 44.77                            & 1.24                            \\
\textit{science}     & 452                              & 53.32                            & 0.90                            \\
\textit{science\_ev} & 125                              & 52.80                            & 0.89                            \\
\textit{food}        & 1555                             & 59.55                            & 0.57                            \\
\textit{science\_wn} & 370                              & 69.73                            & 0.51                            \\ \bottomrule
\end{tabular}
}
\caption{Comparison of \textit{F-score} on original terminology vs filtered terminology. \% Retained refers to the percentage of terms that have single-token hypernyms and are thus retained for the filtered dataset. \% Increase shows the increase in \textit{F-score} on filtered dataset compared to \textit{F-score} on original dataset.}
\label{tab:single-stats}
\end{table}

\section{Conclusion and Future Work}
\label{sec:conclusions}
We have presented a study of different LMs under different settings for zero-shot taxonomy learning. Compared with computationally expensive and highly heuristic methods, our zero-shot alternatives prove remarkably competitive. For the future, we could explore multilingual signals and the integration of traditional word embeddings with contextual representations.

\bibliography{custom}
\bibliographystyle{acl_natbib}

\end{document}